# Evolving Collective Cognition in Human–Agent Hybrid Societies: How Agents Form Stances and Boundaries

### v1

@@
Abstract
Large language models have been widely used to simulate credible human social behaviors. However, it remains unclear whether these models can demonstrate stable capacities for stance formation and identity negotiation in complex interactions, as well as how they respond to human interventions. We propose a computational multi-agent society experiment framework that integrates generative agent-based modeling with virtual ethnographic methods to investigate how group stance differentiation and social boundary formation emerge in human–Agent hybrid societies. Across three studies, we find that agents exhibit endogenous stances, independent of their preset identities, and display distinct tonal preferences and response patterns to different discourse strategies. Furthermore, through language interaction, agents actively dismantle existing identity-based power structures and reconstruct self-organized community boundaries based on these stances. Our findings suggest that preset identities do not rigidly determine the agents’ social structures. For human researchers to effectively intervene in collective cognition, attention must be paid to the endogenous mechanisms and interactional dynamics within the agents’ language networks. These insights provide a theoretical foundation for using generative AI in modeling group social dynamics and studying human–agent collaboration.
In recent years, large language models (LLMs) have made continuous advances in their ability to generate, understand, and respond to natural language, fostering the emergence of a new kind of human–machine sociology. In this emerging context, humans are no longer merely operators of models but co-inhabitants of interactive systems alongside agents capable of social perception and expression [1, 2, 3]. LLM-based are increasingly exhibiting the characteristics of “actors”[4, 5]. The simulation of social behavior has also undergone a fundamental shift, from encoding static rules and actions to the generative exploration of emergent social behavior itself [6]. This transformation gives rise to human–agent Hybrid Societies.
Within such systems, individual cognition is gradually formed through mechanisms such as interaction, imitation, response, and revision. These processes unfold in ongoing language exchanges and social engagement, shaped both by the values of human participants and by the propagation of language among agents [1]. This phenomenon corresponds to what is called group cognition: a dynamic system in which cognition is embedded within community relations, arising through discursive exchanges among group members and irreducible to preset stances held by individuals [7].
From this perspective, agents are not merely tools for task execution but become discursive participants in a shared communicative space. They are capable of forming attitudes, negotiating boundaries, and engaging with institutional structures through interactions with both humans and other agents. While existing studies often focus on the capabilities and controllability of agents [8], they have paid less attention to the generative roles agents may play in constructing social cognition. A key tension arises here: behavioral controllability often depends heavily on preset identities and stances assigned to agents. Many current studies of artificial sociality attempt to simulate society by assigning agents predefined roles and power relations [9, 8]. However, the theory of group cognition suggests that phenomena such as stance differentiation, group affiliation, and boundary negotiation emerge dynamically from interactions among diverse actors rather than from fixed designs [10].
This gap raises a fundamental question: when agents are embedded within social interactions, can they still form attitudes, express stances, and participate in boundary formation through language practices as their preset identities would suggest? Furthermore, can human researchers shape the evolution of collective cognition not through predefinition, but through embodied interventions and co-construction of shared meaning spaces? By building an integrated theoretical and experimental framework, we aim to better understand the formation mechanism of group cognition in a human-agent hybrid society. This understanding points toward a new kind of human–agent sociology [11].
1 The Current Project
Our research focus on the dynamic emergence of stance within agent collectives. Since research on human discourse politics often finds that statistical surveys alone are insufficient to capture the micro-level influence dynamics and the emergence of social cognition within discursive practices [12, 13], and existing multi-agent social simulation frameworks often overlook such capacities, we propose the Computational Multi-Agent Society Experiment (CMASE) framework. This framework introduces the researcher as an embedded participant within a virtual social field, enabling embodied observation, real-time interaction, and qualitative intervention through virtual ethnography method. An advantage of the CMASE framework is that it allows researchers to explore “what-if” scenarios: while the real world permits only a single sequence of events, CMASE enables systematic testing of alternative interventions and their effects on agent behavior.
Based on this framework, we designed a sequence of three studies to explore three central questions: (1) How can human–agent collaborative language interaction be used to influence and shape cognitive convergence or divergence within agent collectives? (2) How are group boundaries of “us” and “them” constructed, negotiated, and disrupted in dynamic interaction? (3) Can preset identities and institutional structures undergo structural transformation through agent participation?
1.1 Study 1
In the first study, we assess the CMASE framework’s ability to replicate real-world sociological mechanisms and evaluate its support for human-embedded participation (Study 1). Specifically, we reconstruct a field study examining the relationship between urban greenery and perceived social fragmentation by replicating its core variables and scenario. By manipulating the level of greenery and enabling the researcher to observe and interact with agents from within the simulated environment, we test the effectiveness of CMASE in simulating complex social settings and generating sociological indicators. We further position the researcher as an embodied actor who intervenes in agent interactions by providing interview and survey-style inputs, thereby investigating whether whether researchers embedded within the scenario can observe finer-grained insights through fieldwork that go beyond the findings of statistical analysis.
1.2 Study 2
The second study investigates the emergence of agents’ “endogenous stances” in multi-agent societies, an underexplored phenomenon that challenges the common assumption that agent attitudes are determined by preset roles. Here, we use the term to refer to stances that deviate from those shaped by the agents’ preset identities and personality traits, instead emerging as reflections of cultural biases embedded in the LLM’s training data [14, 15]. While prior studies have noted such biases, little is known about how they interact with preset identities in shaping collective dynamics.
We aim to explore the possible influence patterns between agents’ preset identities and their endogenous stances and biases, as well as to examine qualitative intervention strategies available to human researchers (Study 2). Specifically, we ask whether LLM-based agents (hereafter referred to as “agents”) in complex disagreement scenarios might exhibit stance tendencies independent of their preset identities. We also examine whether human intervention strategies lose persuasive power when misaligned with these endogenous stances, potentially hindering the reshaping of collective cognition.
We constructed a virtual residential community composed of 30 agents, centered around the contentious issue of siting a waste incineration plant. The agents were assigned to one of three preset groups: environmental advocates, economic supporters, and neutrals, and were given heterogeneous demographic attributes based on real-world survey data. A human researcher was then introduced as a “new resident,” whose discursive interventions combined two stance positions (environmental and economic) with two rhetorical strategies (rational persuasion and emotional provocation), resulting in four experimental conditions. Study 2 is designed to test the following hypotheses:
Hypothesis 1a: agents exhibit endogenous stances independent of their preset identities.
Hypothesis 1b: Alignment between human interventions and agents’ stance tendencies increases the likelihood of influencing collective cognition.
1.3 Study 3
Building on the question raised in Study 2 regarding agents’ stance tendencies beyond preset identities, the third study explores whether such stances can challenge existing authority structures and contribute to the reconstruction of social roles and hierarchies within the community (Study 3). This issue is critical for understanding how emergent cognitive positions might alter the stability and transformation of social structures. Drawing on the intervention strategies examined in Study 2, we conduct a virtual ethnography within CMASE, embedding the researcher into a multi-agent field with preexisting stances and complex social relations. This approach enables close observation of how interactions and events may trigger the dismantling or reconfiguration of social boundaries.
We design a virtual café environment populated by 10 agents. The agents are assigned preset identities, personal goals, and prior opinions about one another to form a preconfigured social hierarchies. Agents in this space interact freely, while the human researcher conducts ethnographic engagement through two approaches: triggering emergency events and participation. Study 3 is designed to test the following hypotheses:
Hypothesis 2a: Social boundaries within the agent collective are not determined by preset identities but emerge gradually through discursive interaction and remain fluid across exchanges.
Hypothesis 2b: Agents are capable of dismantling preset identities through collective action and reconstructing new social structures.
2 Result
2.1 Study 1 validates the CMASE platform’s ability to simulate real-world social dynamics
To evaluate the CMASE platform’s ability to simulate real-world collective behavior, we replicated a recent study in South Korea linking vegetation coverage with community cohesion [16]. This scenario was selected for two main reasons. First, the relationship between vegetation coverage and community cohesion is unlikely to have been directly learned by the LLM, as the original study was published after the release date of the base GPT model used in CMASE. This allows us to assess the platform’s capacity for generating emergent social dynamics rather than recalling known patterns. Second, the phenomenon involves measurable collective outcomes, such as social fragmentation, distrust, exploitation, and indifference, which can be operationalized in a controlled virtual setting and directly compared with empirical human data.

### v2

# Computational Multi-Agents Society Experiments: Social Modeling Framework Based on Generative Agents

@@
Abstract.
This paper introduces CMASE, a framework for Computational Multi-Agent Society Experiments that integrates generative agent-based modeling with virtual ethnographic methods to support researcher embedding, interactive participation, and mechanism-oriented intervention in virtual social environments. By transforming the simulation into a simulated ethnographic field, CMASE shifts the researcher from an external operator to an embedded participant. Specifically, the framework is designed to achieve three core capabilities: (1) enabling real-time human-computer interaction that allows researchers to dynamically embed themselves into the system to characterize complex social intervention processes; (2) reconstructing the generative logic of social phenomena by combining the rigor of computational experiments with the interpretative depth of traditional ethnography; and (3) providing a predictive foundation with causal explanatory power to make forward-looking judgments without sacrificing empirical accuracy. Experimental results show that CMASE can not only simulate complex phenomena, but also generate behavior trajectories consistent with both statistical patterns and mechanistic explanations. These findings demonstrate CMASE’s methodological value for intervention modeling, highlighting its potential to advance interdisciplinary integration in the social sciences. The official code is available at: https://github.com/armihia/CMASE.
LLM-based Agents, Society Simulation, Generative Agent-Based Modeling, Computational Social Science
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Human-centered computing Collaborative and social computingccs: Human-centered computing HCI design and evaluation methods
1. Introduction
The social sciences study human behavior and social structure to understand how society operates. Traditional sociological research relies on human participation to conduct experiments and collect data. Questionnaires and psychological experiments are commonly used to test theoretical hypotheses, understand social phenomena, and predict collective outcomes. Although these methods can provide highly realistic data, they are costly, difficult to scale, and carry certain ethical risks. Therefore, researchers have attempted to solve related problems through agent-based modeling (ABM). However, ABM is mostly applied at a relatively abstract analytical level, which limits its practical utility. It cannot model human behavior and cannot set decision rules for agents. Agents act based on mathematical formulas or logical rules based on parameter values specified by the modeler., which may be influenced by the modeler’s mental model.
Generative agent-based modeling (GABM) offers a novel approach to simulating complex social systems. Vezhnevets et al. (Vezhnevets et al., 2023) pointed out that GABM can capture a richer array of real-world social complexities. In GABM, the modeler does not impose any explicit rules, so the simulation is less biased by the modeler’s mental framework. Ghaffarzadegan et al. (Ghaffarzadegan et al., 2024) further note that GABM allows modelers to define distinct personality traits for each agent, including age, gender, temperament, occupation, and other characteristics.
Xue et al. (Xue et al., 2024) propose a two-dimensional classification of sociological modeling based on (1) whether there is human intervention and (2) whether the model focuses on feature explanation or outcome prediction. This yields four types: descriptive modeling (no intervention, focus on specific features or effects), interpretative modeling (intervention, focus on specific features or effects), predictive modeling (no intervention, focus on forecasting outcomes), and comprehensive modeling (intervention, focus on forecasting outcomes). ABM and its derivative GABM belong to descriptive modeling, treating social simulation as a virtual laboratory where researchers can explore various social phenomena. Predictive modeling, by contrast, focuses on outcome forecasting and often integrates ABM with other techniques (such as time-series analysis, forecasting tournaments, and supervised machine learning) to understand and predict the behavior of complex social systems.
Although traditional and computational social science research (Lazer et al., 2009) has extensively addressed descriptive modeling, interpretative modeling, and predictive modeling, comprehensive modeling, which involves human intervention and focuses on forecasting outcomes, remains relatively unexplored. This approach requires models not only to simulate the dynamic behavior of complex social systems, but also to incorporate and respond to human interventions. These interventions may take the form of policy changes, institutional adjustments, or individual behavioral inputs. While conventional modeling typically simplifies intervention as fixed parameter settings, comprehensive modeling emphasizes the effects of the intervention process itself. Models of this kind must capture the system’s initial state before the intervention, dynamically trace its evolution during intervention, and ultimately produce predictions about the likely consequences afterward. Therefore, comprehensive modeling is not merely concerned with explaining social mechanisms, but is intended to support forward-looking assessments and real-world decision-making.
While recent advancements have yielded powerful large language model (LLM)-based multi-agent platforms to operationalize GABM, most of these systems remain constrained to descriptive or predictive tasks. Despite excelling at scaling up agent populations or simulating specific static scenarios, they typically function as closed environments where human involvement is limited to the initial setup and post-hoc data analysis. They lack the interactive mechanisms necessary for researchers to dynamically embed themselves into the ongoing simulation. This highlights a critical gap: the absence of a framework that supports real-time, human-in-the-loop intervention to organically observe, steer, and interpret complex social dynamics.
In summary, to bridge the limitations of existing sociological modeling methods, this study proposes a comprehensive modeling framework to achieve the following research objectives:
  • To enable real-time human-computer interaction (HCI) within social simulations. This objective replaces the static, predefined rules of traditional agent-based modeling with a dynamic environment. By embedding researchers directly into the simulation loop, the framework allows for the continuous observation and adjustment of complex social processes as they unfold.
  • To implement and evaluate researcher-led interventions. Rather than treating simulation as a black box, this objective utilizes an ethnography-inspired approach to conduct controlled interventions. This allows researchers to test how specific changes in variables or agent interactions influence the generative logic of social phenomena.
  • To provide predictive insights grounded in causal explanation. This objective focuses on using simulation data to forecast the outcomes of real-world policy or social changes. Unlike purely data-driven models that identify correlations, this framework aims to demonstrate the underlying mechanisms of intervention effects, providing a causal basis for decision-making.
Based on this understanding, the present study draws on insights from GABM (Vezhnevets et al., 2023) and computational experiments (Xue et al., 2024) to develop a novel framework for comprehensive modeling, referred to as Computational Multi-Agents Society Experiments (CMASE). This framework not only builds on multi-agent simulation but also introduces a real-time mechanism for human intervention. It allows researchers to observe, adjust, and interact with the model environment as it runs, much like participating in a real-world field study.
CMASE transforms the simulation environment into a kind of simulated ethnographic field, where the researcher is no longer an external operator of a static model but becomes an embedded participant or social actor within the system. This approach bridges the gap between computational modeling and field-based research, making the model itself a new venue for conducting computational ethnography.
2. Related Work
2.1. Simulating Social Behavior with Large Language Models
Social simulation has been an important tool for understanding and analyzing social phenomena. Early studies mainly relied on rule-based artificial agent models, namely agent-based modeling (ABM)(Bonabeau, 2002; Macy and Willer, 2002). Traditional ABMs often depend on manually crafted heuristic rules, making it difficult to approximate the complex features of real human cognition, language, and emotion (Ma et al., 2024; Wu et al., 2023; Jäger and Reisinger, 2022). In recent years, with the capabilities of Large Language Models (LLMs) in simulating human behavior, research on building social simulators based on LLMs has rapidly advanced (Park et al., 2023). GABM has gradually become one of the core methods for simulating complex social behavior (Vezhnevets et al., 2023). In this individual-level approach, mechanistic models of agent interaction are combined with LLMs, and each agent makes decisions after communicating with the LLM. This innovative method minimizes the need for modelers to make assumptions about human decision-making, instead leveraging the vast data embedded in LLMs to capture human behaviors and choices (Ghaffarzadegan et al., 2024).
Existing GABM systems can be roughly categorized into two main research directions. The first line of research focuses on realistic simulations of specific social scenarios, aiming to foster competition or collaboration among agents to generate social insights or support policy design. Li et al. (Li et al., 2024) developed Agent Hospital, where LLM-driven doctor agents interact with large numbers of patients to evolve transferable medical decision-making capabilities. Yu et al. (Yu et al., 2024) designed FinCon, a hierarchical agent system modeled after financial institutions, using language interaction and belief reinforcement to enhance collective collaboration in complex investment tasks. TrendSim (Zhang et al., 2024a) focuses on topic propagation on social media and its vulnerability to poisoning attacks, providing intervention strategies for platform governance. ElectionSim (Zhang et al., 2024b) reconstructs election scenarios using millions of social media records, achieving high-fidelity voter modeling and policy preference prediction. Zhao et al. (Zhao et al., 2023) simulated competitive dynamics between restaurant agents and customer agents to align with existing market and sociological theories.
The second line of GABM research emphasizes building reusable, scalable simulation platforms that support general modeling tasks. Wang et al. (Wang et al., 2025) divide the development of LLM-based social simulation platforms into two stages. In the first stage, researchers focus on simulating specific scenarios with typically fewer than 1,000 agents (such as (Zhou et al., 2025)). By proposing YuLan-OneSim, they aim to push the field into the next stage by introducing environmental construction, agent evolution, and feedback integration mechanisms, thus enabling simulations of larger and more complex social processes. Frameworks of similar scale now support simulations with over 10,000 agents. AgentSociety (Piao et al., 2025) recreates contexts to evaluate policy interventions and the emergence of social mechanisms. AgentScope (Gao et al., 2024) offers modular design, no-code interfaces, and distributed deployment, significantly lowering the barrier to building multi-agent systems. Socioverse (Zhang et al., 2025) integrates data from tens of millions of real users to provide a reliable foundation for modeling multiple subdomains such as politics and economics. GenSim (Tang et al., 2024) supports the operation of up to 100,000 agents and includes error correction mechanisms, emphasizing generalizability and long-term robustness. OASIS (Yang et al., 2024) uses open social platforms as simulation environments, supporting dynamic networks, recommendation systems, and multi-behavior modeling.
However, exponential increases in agent quantity alone are not sufficient to effectively capture the essence of social complexity. Recent research on freely formed AI collectives similarly suggests that social diversity, innovation, and self-regulation emerge less from scale than from the dynamics of open-ended interaction (Shiyang et al., 2024). In particular, without mechanisms for understanding social context and interactive feedback, large-scale simulations often struggle to offer human-interpretable responses to core questions such as the formation of social structures, norms, or institutional evolution.
Unlike early studies that relied on extensive handcrafted heuristic rules to dictate specific actions, our approach only establishes a priori environmental constraints and cognitive workflows. Within these boundaries, the system relies entirely on the LLM-based agents’ capacity to replicate human behavior to drive the simulation. This ensures that the resulting social dynamics are systematic, emergent properties rather than cherry-picked, pre-defined outcomes.
2.2. Human-AI Collaborative Modeling in Social Systems
Computational social science explores social structures, behavioral mechanisms, and institutional evolution in simulated computational environments through formalized approaches (Lazer et al., 2009). Within this field, the computational experiments paradigm aims to systematically manipulate variables in virtual societies, thereby extending the boundaries of traditional experimentation (Xue et al., 2023). In this context, LLM-based GABM is not only a representational tool for social processes but also holds the potential to become a platform for computational experimentation. Sato (Sato, 2024) highlights the importance of meaning and reflexivity in bridging the gap between agent-based modeling and social theory. This implies that for agent-based modeling to contribute to theoretical advancement in the social sciences, it must go beyond scale amplification and incorporate workflows that offer human interpretability. Therefore, beyond the expansion of scale, a more socially grounded, interactive, and reflexive human–AI sociological framework is needed, in which agents are no longer passive generators of behavior but are embedded within social networks and actively participate in meaning-making processes (Tsvetkova et al., 2024).
Recent research in this direction has focused on the subfield of human-AI co-creation. Broadly speaking, GABM belongs to this category: through human-designed environments, agent adaptation, and human observational feedback, it enables collaborative modeling between humans and AI. Some studies emphasize interactive environment co-construction. For example, HABITAT 3.0 (Puig et al., 2023) supports real-time task collaboration between humans and humanoid agents in virtual spaces. As this field has progressed, researchers have begun to explore more creative and socially complex scenarios of human-AI co-creation, emphasizing how humans guide emergent AI behavior. The Cognitio Emergens framework (Lin, 2025) redefines human-AI collaboration as a process of mutual evolution and offers a conceptual toolkit to preserve the significance of human participation. The Pecan method (Lou et al., 2023) enhances zero-shot collaboration between AI and humans via strategic integration and context-aware mechanisms. The Collaborative Gym framework (Shao et al., 2024) highlights the human role in environmental control and structural intervention during task execution, supporting asynchronous three-way interaction between humans, agents, and task environments.
3. CMASE Framework
Computational social science is an emerging interdisciplinary field that leverages digital technologies, large-scale data analysis, and computational methods to study social phenomena (Conte et al., 2012). Xue et al. (Xue et al., 2024) propose a two-dimensional classification of Computational sociological modeling based on (1) whether there is human intervention and (2) whether the model focuses on feature explanation or outcome prediction. This yields four types: descriptive modeling (no intervention, focus on specific features or effects), interpretative modeling (intervention, focus on specific features or effects), predictive modeling (no intervention, focus on forecasting outcomes), and comprehensive modeling (intervention, focus on forecasting outcomes). Generative agent-based modeling (GABM) belong to descriptive modeling, treating social simulation as a virtual laboratory where researchers can explore various social phenomena. Vezhnevets et al. (Vezhnevets et al., 2023) pointed out that GABM can capture a richer array of real-world social complexities. In GABM, the modeler does not impose any explicit rules, so the simulation is less biased by the modeler’s mental framework. Ghaffarzadegan et al. (Ghaffarzadegan et al., 2024) further note that GABM allows modelers to define distinct personality traits for each agent, including age, gender, temperament, occupation, and other characteristics.
Although traditional and computational social science research (Lazer et al., 2009) has extensively addressed descriptive modeling, interpretative modeling, and predictive modeling, comprehensive modeling, which involves human intervention and focuses on forecasting outcomes, remains relatively unexplored. This approach requires models not only to simulate the dynamic behavior of complex social systems, but also to incorporate and respond to human interventions. These interventions may take the form of policy changes, institutional adjustments, or individual behavioral inputs. While conventional modeling typically simplifies intervention as fixed parameter settings, comprehensive modeling emphasizes the effects of the intervention process itself. Models of this kind must capture the system’s initial state before the intervention, dynamically trace its evolution during intervention, and ultimately produce predictions about the likely consequences afterward. Therefore, comprehensive modeling is not merely concerned with explaining social mechanisms, but is intended to support forward-looking assessments and real-world decision-making.
Based on this understanding, the present study draws on insights from GABM (Vezhnevets et al., 2023) and computational experiments (Xue et al., 2024) to develop a novel framework for comprehensive modeling, referred to as Computational Multi-Agents Society Experiments (CMASE). Drawing on natural language descriptions similar to those used in tabletop role-playing game (TRPGs, a game where players use only dialogue and dice to drive imagined adventures) (White et al., 2024), this framework supports the creation of highly customizable social simulation environments. It facilitates the exploration of how simple interactions among multiple agents give rise to complex collective behaviors or patterns, thereby replicating the intricate dynamics of real-world societies. At the same time, it incorporates human intervention by embedding human researchers within the agent environment to actively shape the simulation. In addition, unlike descriptive modeling approaches such as GABM, CMASE transforms the simulation environment into a kind of simulated ethnographic field, where the researcher is no longer an external operator of a static model but becomes an embedded participant or social actor within the system. This approach bridges the gap between computational modeling and field-based research, making the model itself a new venue for conducting computational ethnography. This allows them to actively guide specific directions of social evolution through live interaction. As shown in Figure 1, the overall structure consists of four components: environment maker, environment, agent, and event. To facilitate practical application and reproducibility, we provide the complete source code, detailed setup instructions, and a step-by-step example for initializing new experiments in our open-source repository: https://github.com/armihia/CMASE
Figure 1. Overall Structure of CMASE.
3.1. Environment Maker
Using textures, creators can define areas on the map. There are two types of areas: region areas and object areas. Each area consists of a collection of textures. A region area is a set of ground textures and includes a description of that region. An object area is a set of furniture or item textures. To support realistic social simulations, the CMASE framework provides an interface for creators to assign functional attributes that align with daily usage scenarios. Specifically, creators can define how an object alters an agent’s state based on its real-world utility. For example, reflecting common daily routines, a creator might configure a chair to add a “feeling comfortable” state to an agent, or a repair tool to remove a “damaged” state from a table.
3.2. Environment
The environment module uses the map created by the environment maker as its base. Based on this foundation, researchers can define the distribution of demographic variables for agents to batch-generate generative agents that match the research sample. Researchers may also assign distinct long-term goals (e.g., “maximize social connections within the community,” “accumulate specific resources,” or “maintaining a low profile”) and initial items to agents according to their demographic characteristics. These goals serve as persistent motivational drivers influencing agent behavior throughout the simulation.
Additionally, the module can provide agents with perceptual information and action conflict detection in both image and text form, and it can visualize environmental data as statistical charts. At the start of each round, the environment calculates the perceptual information for each agent based on their current grid location and line-of-sight obstructions (e.g., walls). It then synthesizes visible environmental data—including nearby terrain, interactive objects, the presence and visible actions of other agents, and recent local events—into structured textual descriptions (and optional visual representations). This synthesized data forms the contextual prompt provided to the agent’s cognitive module for decision-making.
Time in the environment advances in rounds. To ensure balanced agent participation and maintain manageable simulation steps, we adopted constraints inspired by traditional tabletop role-playing game (TRPG) systems, including Dungeons & Dragons (D&D). These systems provide time-tested heuristics for discretizing continuous human actions into manageable rounds. Consequently, one round corresponds to approximately fifteen seconds in the real world, that is a standard duration in tactical simulations sufficient for completing a single, meaningful interaction. Within a single round, each agent may perform one standard action and one movement action. Movement is limited to a maximum of twenty grid cells per round to mimic realistic movement speeds (approx. 30-60 ft) relative to the map scale. If the standard action is “chat with [agent]” (see section 3.3.4), the chat content is limited to thirty words to simulate concise, real-time exchanges and prevent excessive dialogue generation in a single turn.
3.3. Agent
3.3.1. Affective Variables
To ensure that agents in CMASE can closely simulate individuals with agency in computational social systems, a new proactive architecture incorporating emotion and desire is needed. In this framework, ’needs’ are defined as the agent’s internal motivational states, distinct from external demands. These needs are concretized into long-term and short-term goals. Long-term goals are defined by the user during agent generation based on the research objectives and remain unchanged afterward. Short-term goals are automatically updated in each round as the agent generates its next action.
In psychological research, Ekman’s basic emotion model has been criticized for its coarse categorization and static view of emotions. Recent research draws on the phenomenological concept of “intentionality” to propose a restructured model of emotion. This model represents emotional states as principal component vectors rather than discrete outward expressions, such as happiness, and treats emotion as an internal representation shaped by external stimuli through self-reflection. Based on this perspective, the emotions of agents in CMASE are modeled using the VAD emotion framework (Russell and Mehrabian, 1977).
Specifically, we utilize the NRC VAD Lexicon v2 (Mohammad, 2025) to map the textual description of the agent’s short-term situational cognition to a quantitative emotional vector. The process involves two steps. First, the input text is preprocessed using lemmatization and maximum matching to align words and phrases with the lexicon. The agent’s current state vector $\mathbf{E}_{t}=[v,a,d]$ is then calculated as the arithmetic mean of the Valence, Arousal, and Dominance scores of all identified lexical units in the text.Second, to bridge the gap between continuous numerical vectors and the LLM’s linguistic reasoning, we employ a threshold-based discretization method. The continuous range of each dimension is divided into five equidistant intervals, mapping the numerical scores to natural language descriptors: Low, Relatively Low, Medium, Relatively High, and High. These descriptors are explicitly embedded into the agent’s prompt to dynamically constrain and modulate its subsequent reasoning and action generation.
3.3.2. Cognitive Variables
Cognitive variables include an agent’s demographic variables, long-term self-awareness, and short-term situational cognition. Demographic variables are sampled directly from the overall distribution provided by the environment and are used during initialization to generate the agent’s long-term self-awareness. These two long-term properties do not change after initialization to preserve the consistency of the agent’s personality.
In each round, when the agent makes a decision, the LLM generates new short-term situational cognition based on the agent’s long-term self-awareness and the current context. This cognition is analyzed to produce VAD emotional values, which are then converted into a natural language description. This description is fed back to the agent to constrain the direction of its subsequent actions. Essentially, the emotional description dictates the logical boundaries for the agent’s behavior. For example, if the VAD values describe a ’furious’ state, the agent is constrained to generate impulsive actions rather than rational responses. In addition, when an agent’s state changes due to actions or events, its state is appended to the short-term situational cognition so that it can be perceived by the agent itself.
3.3.3. Memory Variables
3.3.4. Action Space
After an agent receives information from perceptual inputs and its internal attributes, it structures that information and passes it to the LLM to determine the next action. There are seven action types available: (1) “go to [area],” a movement action that employs the $A^{*}$ algorithm to automatically compute an optimal path to a target location within the agent’s perceptual range; (2) “use [furniture, item],” which causes the agent to employ the specified object; (3) “apply [item] to [agent, furniture],” by which the agent uses an item to interact with another agent or furniture; (4) “take [item],” enabling the agent to acquire an object; (5) “put [item] in/on [area, furniture],” allowing the agent to place an object into or onto a designated location or piece of furniture; (6) “give [item] to [agent],” through which the agent transfers a held object to another agent; and (7) “chat with [agent],” whereby the agent initiates a dialogue with one or more peers, with the content of the dialogue generated at decision time.
3.4. Event
Emergency events are explicitly designed by researchers to mirror real-world crises and specific social scenarios required for the simulation. When predefined conditions are met, these pre-configured events are activated to change the state of objects within a specified range, defined either by coordinates or by an area. Emergency events are classified as either scheduled or triggered. Scheduled emergency events activate automatically after a set number of rounds. Triggered emergency events activate in two ways: an existence trigger fires when the environment detects an object of a certain type or description, and an action trigger fires when an agent performs a certain action.
4. Platform Demonstration: Simulating Context-Dependent Social Dynamics
To evaluate the CMASE platform’s ability to simulate real-world collective behavior, we reconstruct a field study examining the relationship between urban greenery and perceived social fragmentation by replicating its core variables and scenario (Lee and Han, 2025). This scenario was selected for two main reasons. First, while we acknowledge that the general positive correlation between green spaces and social cohesion is well-documented in prior literature likely included in the model’s pre-training data (e.g., Wan et al. (Wan et al., 2021)), the specific multidimensional framework of social fragmentation (distrust, exploitation, and indifference) proposed by Lee and Han (Lee and Han, 2025) offers a distinct validation target. By replicating these specific psychological mechanisms rather than just general trends, we aim to assess the platform’s capacity for generating emergent social dynamics. Second, beyond mere validation, this study aims to extend the original data-driven research by establishing a real-time generative field that enables dynamic inquiries, such as agent interviews and virtual field observations, which were not possible in the original static study.
4.1. Setup
We selected a study published in a top journal and replicated its survey results under comparable conditions. To ensure that the LLM had not already been trained on the studies being replicated, while acknowledging that this measure cannot strictly guarantee the complete exclusion of broad theoretical trends from prior works, we employed the GPT-4o “2024-12-01-preview” model and limited our replication to the paper published after that date. Crucially, beyond validation, this study aims to extend the original data-driven research by establishing a real-time generative field that enables dynamic inquiries, such as agent interviews and virtual field observations.
Lee and Han (Lee and Han, 2025) analyzed interview data collected in South Korea during the COVID-19 pandemic and demonstrated that regions with higher vegetation cover exhibited lower levels of social fragmentation, indicating stronger community bonds. Based on the descriptions in the original study, we used the environment maker module to construct maps of South Korean regions with varying degrees of vegetation cover. Because the original paper emphasized South Korea’s community isolation policies during the pandemic, we designed an isolated community environment and placed 10 agents in it.
Specifically, the simulated community comprises ten individual residential units surrounding a shared public green space. To support diverse activities, the green space was designed with distinct functional zones, including a recreation area, a rest zone, and a scenic lakeside area equipped with benches. We placed ten agents in this environment, each initialized with a distinct persona, including a specific name, age, and occupation (e.g., lawyer, fashion designer, software engineer), to simulate a heterogeneous micro-society.
4.2. Measures